OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8

rishavraj (rishavraj):

from first equation x = 6 + 2y

rishavraj (rishavraj):

on plugging the value of x in second equation \[3(6 + 2y) + y = 8 \]

OpenStudy (anonymous):

i forgot to put answers here they are -2y + 6 2y - 6 2y + 6 -2y - 6

OpenStudy (owlcoffee):

All you have to do, is take the expression that Rish wrote above and simplify it as much as you can.

OpenStudy (anonymous):

is that JB in your picture and im not good at simplifying equations like this lolol

OpenStudy (owlcoffee):

No, that's my own photo... And I think Rish already gave you the answer.

OpenStudy (anonymous):

2y+6?

rishavraj (rishavraj):

@TheAlgebraHelper man i wrote the answer in my very first comment.....:)

OpenStudy (anonymous):

so it is 2y+6 LOL

OpenStudy (anonymous):

@rishavraj help me with one more?

rishavraj (rishavraj):

okay....

OpenStudy (anonymous):

An equation is shown below: 3 over 2 multiplied by x plus 7 over 2 equals 2 to the power of x What is the solution to the equation?

OpenStudy (anonymous):

x = 1 x = 3 x = 5 x = 8

rishavraj (rishavraj):

u mean \[\frac{ 3 }{ \frac{ 2x + 7 }{ 2 } } = 2^{x}\]

OpenStudy (anonymous):

umm no

OpenStudy (anonymous):

\[3\over2\] x +\[7\over2\] = 2^x

rishavraj (rishavraj):

okay then x = 3......

OpenStudy (anonymous):

lol thnx